OpenStudy (anonymous):

Juan buys 4 shirts on sale for $15 each. He also buys 6 pairs of socks, each pair of socks being the same price. If the total cost for the clothes is $78, how much did one pair of socks cost? A. $3 B. $8 C. $3.70 D. $3.50

OpenStudy (anonymous):

If the total cost for the clothes is $78, then one pair of socks cost $3.00. ...4 shirts @15.00 each = 60.00. [ $78.00-60.00 = 18.00 18/6=$3.00 ] So A.

OpenStudy (johnweldon1993):

Well...we know the total cost is 78 And we know he spent (15 * 4) = 60 dollars on shirts... so now he has 78 - 60 = 18 dollars to work with... He wants to buy 6 socks...at the same price...to make that 18 dollars...so 6x = 18 divide both sides by 6 and solve for 'x'